OpenStudy (anonymous):

Assume that there are 6 different issues of Sports Illustrated, 7 different issues of Time, and 3 different issues of Newsweek, including the December 1st issue, on the rack. You choose 4 of them at random. (1) What is the probability that exactly 1 is an issue issue of Sports Illustrated? (2) What is the probability that you choose the December 1st issue of Newsweek?

Directrix (directrix):

(1) [C(6,1) C(10,3)] / C(16,4) = 36/91 = .396 approx

Directrix (directrix):

(2) [C(1,1) C(15,3)] / C(16,4) = 1/4 = .25